Exercise 2-4 (Video)

Manufacturing cost data for Orlando Company, which uses a job order cost system, are presented below.

Indicate the missing amount for each letter. Assume that in all cases manufacturing overhead is applied on the basis of direct labor cost and the rate is the same. (Round overhead rate to 2 decimal places, e.g. 15.25 and final answers to 0 decimal places, e.g. 5,275.)

Case A

Case B

Direct materials used $

(a) $90,100
Direct labor 60,000 146,100
Manufacturing overhead applied 39,000

(d)
Total manufacturing costs 154,750

(e)
Work in process 1/1/20

(b) 16,600
Total cost of work in process 207,900

(f)
Work in process 12/31/20

(c) 12,100
Cost of goods manufactured 198,200

(g)
Answer #1

Case A

Total manufacturing cost = Direct materials + Direct labor + Manufacturing overhead applied

154,750 = Direct materials + 60,000+39,000

Direct materials = $55,750

Total cost of work in process = Total manufacturing cost +Work in process, 1/1/20

207,900 = 154,750 + Work in process, 1/1/20

Work in process, 1/1/20 = $53,150

Cost of goods manufactured = Total cost of work in process - Work in process 12/31/20

198,200 = 207,900- Work in process 12/31/20

Work in process 12/31/20 = $9,700

Manufacturing overhead rate in Case A = Manufacturing overhead applied / Direct labor

= 39,000/60,000

= 65%

Case B

Manufacturing overhead applied = Direct labor x Manufacturing overhead rate

= 146,100 x 65%

= $94,965

Total manufacturing cost = Direct materials + Direct labor + Manufacturing overhead applied

= 90,100+146,100+94,965

= $331,165

Total cost of work in process = Total manufacturing cost +Work in process, 1/1/20

= 331,165+16,600

= $347,765

Cost of goods manufactured = Total cost of work in process - Work in process 12/31/20

= 347,765-12,100

= $335,665

Case A Case B
Direct materials used //img.homeworklib.com/questions/babd6d00-77b5-11ea-85a4-975cf49c3d52.gif?x-oss-process=image/resize,w_560
55,750
(a) $90,100
Direct labor 60,000 146,100
Manufacturing overhead applied 39,000 //img.homeworklib.com/questions/babd6d00-77b5-11ea-85a4-975cf49c3d52.gif?x-oss-process=image/resize,w_560
94,965
(d)
Total manufacturing costs 154,750 //img.homeworklib.com/questions/babd6d00-77b5-11ea-85a4-975cf49c3d52.gif?x-oss-process=image/resize,w_560
$331,165
(e)
Work in process 1/1/20 //img.homeworklib.com/questions/babd6d00-77b5-11ea-85a4-975cf49c3d52.gif?x-oss-process=image/resize,w_560
53,150
(b) 16,600
Total cost of work in process 207,900 //img.homeworklib.com/questions/babd6d00-77b5-11ea-85a4-975cf49c3d52.gif?x-oss-process=image/resize,w_560
$347,765
(f)
Work in process 12/31/20 //img.homeworklib.com/questions/babd6d00-77b5-11ea-85a4-975cf49c3d52.gif?x-oss-process=image/resize,w_560
9,700
(c) 12,100
Cost of goods manufactured //img.homeworklib.com/questions/babd6d00-77b5-11ea-85a4-975cf49c3d52.gif?x-oss-process=image/resize,w_560
198,200
//img.homeworklib.com/questions/babd6d00-77b5-11ea-85a4-975cf49c3d52.gif?x-oss-process=image/resize,w_560
$335,665
(g)

Kindly give a positive rating if you are satisfied with this solution and please ask if you have any query.

Thanks